OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2). y = 2x + 4 or y = − 1/2x − 7/2 but i feel like i did something wrong so are they both wrong?

OpenStudy (anonymous):

parallel to means same slope as the line given. use y=mx+b equation and substitute things. (x,y) and the m (slope) to find b. If it isn't obvious enough, (3,-2) is the (x,y)

OpenStudy (anonymous):

but the 2 and 4 are in for the slope and y. so this is where I'm kinda stuck.

OpenStudy (anonymous):

wait nvm. think i got it.

OpenStudy (anonymous):

i got -5. but my options are. y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8